# Accelerated Bayesian Inference for Pulsar Timing Arrays: Normalizing Flows for Rapid Model Comparison Across Stochastic Gravitational-Wave Background Sources

### v1

@@ VI Summary
In this work, we present a normalizing-flow-based machine learning (NF-based ML) framework for stochastic gravitational-wave background (SGWB) model selection using pulsar timing array data – the first application of ML to SGWB model comparison. In our approach, conditional normalizing flow networks were trained on the NANOGrav 15-year dataset and incorporated a learned harmonic mean estimator to directly infer Bayesian posteriors and model evidences (Bayes factors). We tested ten representative SGWB source models spanning both astrophysical and cosmological scenarios. Despite the high dimensionality (22 parameters per model), the normalized flow-based inference completes in only $\sim$20 hours per model, compared to roughly $\sim$ 10 days for MCMC analyses.
The posterior distributions obtained with the normalizing flows are in good agreement with those from traditional MCMC sampling, with Hellinger distances typically $\lesssim 0.3$ (on a 0–1 scale where 0 indicates identical distributions). Likewise, the Bayes factors derived from the NF-based ML framework agree with MCMC-based calculations within their reported uncertainties, correctly ranking the evidence for each SGWB model. These findings demonstrate that our ML-driven approach achieves comparable accuracy to standard Bayesian inference while reducing runtime by an order of magnitude or more. In summary, this work provides a robust and faster framework for SGWB model selection, one that is immediately applicable to current PTA datasets and well suited for the demanding analyses of near-future PTA data.
@@ Appendix A Data
Pulsar Timing Residuals White Noise [ns]
midrule J0030+0451 724 685.7

### v2

@@ VI Summary
In this work, we present a normalizing-flow-based machine learning (NF-based ML) framework for stochastic gravitational-wave background (SGWB) model selection using pulsar timing array data – the first application of ML to SGWB model comparison. In our approach, conditional normalizing flow networks were trained on the NANOGrav 15-year dataset and incorporated a learned harmonic mean estimator to directly infer Bayesian posteriors and model evidences (Bayes factors). We tested ten representative SGWB source models spanning both astrophysical and cosmological scenarios. Despite the high dimensionality (22 parameters per model), the normalized flow-based inference completes in only $\sim$20 hours per model (10 pulsars), compared to roughly $\sim$ 10 days for MCMC analyses (68 pulsars). Note that the substantial time reduction is partly due to the decrease in dataset size from 68 pulsars to 10 pulsars. It may be also partly attributable to the use of different computers—NF‑ML was run on a GPU computer, while MCMC used a CPU computer—even though the GPU computer for NF‑ML is roughly three times less expensive (See Sec. F for hardware specifications). In the future, more pulsars will be taken into account. To test scalability, we have run epochs for $N_{\mathrm{pulsars}}=(8,9,10,11,12,13,14,15)$ with corresponding numbers of time residuals $N_{\mathrm{time-residual}}=(3797,4220,4944,$ 5604,6228,6585,7063,7463)5604,6228,6585,7063,7463)5604 , 6228 , 6585 , 7063 , 7463 ) with IGW model (Eq. (31)). Fitting these points (Fig. 10, based on Table 4) yields an average time per residual per epoch of
$T_{\mathrm{per-res}}\simeq 0.13\,\mathrm{s},$ (17)
which corresponds to the slope of the fitted line. These results suggest that per‑pulsar processing time—assuming a similar number of residuals per pulsar—remains roughly constant, and that total training time grows nearly linearly. However, we acknowledge that measurements at small $N_{\mathrm{pulsars}}$ are insufficient to confirm this trend at larger scales. Due to computational resource limits, a full test at $N_{\mathrm{pulsars}}=68$ (with $N_{\mathrm{time-residual}}=20290$) is not yet feasible, even though the average number of time residuals per pulsar decrease from about 500 to 300. We therefore describe our scalability conjecture as preliminary and plan to extend these studies when additional resources become available. [^3] We appreciate the anonymous Referee for highlighting this point to us.
Figure 10: Relationship between training time per epoch and the number of time residuals. The slope of $0.13$ seconds reflects the average time per residual per epoch, $T_{\mathrm{per-res}}$, while the intercept of $224.97$ seconds captures time for file loading, model initialization, and saving.
$N_{\mathrm{pulsars}}$ $N_{\mathrm{time-residual}}$ Training time per epoch (Seconds)
8 3797 694
9 4220 775
10 4944 873
11 5604 920
12 6228 989
13 6585 1033
14 7063 1106
15 7463 1207
Table 4: Summary of Pulsar Data Used in the Scalability Analysis.
The posterior distributions obtained with the normalizing flows are in good agreement with those from traditional MCMC sampling, with Hellinger distances typically $\lesssim 0.3$ (on a 0–1 scale where 0 indicates identical distributions). Likewise, the Bayes factors derived from the NF-based ML framework agree with MCMC-based calculations within their reported uncertainties, correctly ranking the evidence for each SGWB model. These findings demonstrate that our ML-driven approach achieves comparable accuracy to standard Bayesian inference while it may significantly reduce runtime by using a GPU setup. In summary, this work provides a robust and faster framework for SGWB model selection, one that is immediately applicable to current PTA datasets and well suited for the demanding analyses of near-future PTA data.
@@ Appendix A Data
Name Timing Residuals White Noise [ns]
J0030+0451 724 685.7